# Topological photon pumping in quantum optical systems

### v2

@@ 4.2.1 Photon pumping on a two-level atom chain
Figure 7: Photon pumping on a waveguide system. a: Dependence of the extended hopping rates $\bar{J}^{\prime}$ and $\bar{J}$ with the angle $\phi^{\prime}_{1}$ (all rates are normalized to the maximal rate $\bar{J}_{\text{max}}\equiv|\text{max}\{\bar{J}(t)\}|$ during the cycle). The gray shaded area represents $\delta\bar{J}<0$, where the winding number of the extended SSH model is $\nu=1$ (TP: topological phase). b: Variation of the angle $\phi^{\prime}_{1}(t)$ during the pumping cycle. Importantly, twice during the cycle the extended hopping rates coincide, such that $\delta\bar{J}=0$. c: We vary the energy offset as $\Delta(t)=\Delta_{\mathrm{max}}\sin(2\pi t/T+\pi/2)$ such that the path in parameter space encircles the topological degeneracy point ensuring the displacement of the center of mass after a completed cycle. d: The left column shows the time-integrated Berry curvature of the lower band ${\cal W}_{k}^{-}$ and quasi-momentum distribution density $|f(k)|^{2}$ (in arbitrary units). The right column shows the spatial probability density of the wave packet in real space (on the $B$ sublattice), initially (dashed line) and after its numerical evolution for ten pumping cycles (solid line). The width of the Gaussian wave packet in momentum space $w_{k}$ (the same as in Figs. 5 and 6) is doubled from the upper to the lower row, leaving the displacement of the photon’s center of mass $\Delta x=10\cdot 0.919a$ unchanged, but leading to an increase of the dispersion, such that the fidelity per cycle ${\cal F}(T)=0.998$ and after ten cycles ${\cal F}(10T)=0.962$ decreases to 0.987 and 0.540, respectively.
@@ 4.3 All-to-all hopping: Waveguide system
Finally, we discuss a chain of emitters coupled to a waveguide. This setup can be realized, for instance, with cold atoms trapped in the vicinity of a tapered optical fiber [57, 74, 75]. In this system, the coherent interactions between atoms are mediated by the fundamental guided modes of the waveguide. Note, that we limit the discussion in this paper to cases where the coupling efficiency of the emitters to the guided modes is almost $100\%$, as it can be achieved, for example, with artificial atoms in semiconductors (quantum dots) [76, 77]. In general, however, one needs to consider the coupling through the radiation or unguided modes. In the case of atoms placed around a cylindrical nanofiber (see Fig. 4d), the interaction between atoms separated by an angle $\phi_{ij}=\phi_{j}-\phi_{i}$ and a distance $z_{ij}=z_{j}-z_{i}$ are given by
@@
where $\gamma=\frac{2|d|^{2}\omega_{0}\beta^{\prime}}{\hbar\epsilon_{0}}|e_{r}|^{2}$ is the rate of a photon to be emitted into the waveguide for a single atom. Moreover, we consider in particular the situation where the atomic cells are positioned as a helix around the waveguide as shown in Fig. 4d. The hopping rates in this case are given by
@@
where $\phi_{1}$ and $\phi_{1}^{\prime}$ are the angle between atoms in neighboring cells of the same sublattice, and between atoms of different sublattices within the same cell, respectively. In order to ensure that the chiral symmetry is not broken, we fix the value of $a$ to $a=\pi/\beta$. We also fix the values of $b=a/2=\pi/(2\beta)$ and $\phi_{1}=\pi/2$, such that $J_{2p}=0$ and the remaining hopping rates read
@@ 4.3.1 Photon pumping on a waveguide system
As shown in Fig. 7a, the transition from a topological to a non-topological phase occurs in this system when the angle $\phi_{1}^{\prime}=\pi/4$, since at this point $J_{2p-1}=J^{\prime}_{2p-1}$ for all values of $p$. A pumping scheme can then be established by adiabatically rotating sublattice $B$ by an angle $\phi_{1}^{\prime}(t)$ which after a time $T$ returns to the initial configuration (Figs. 7b and c) [78]. As it was discussed for atoms in free space (sec Sec. 4.2), due to the long-ranged character of the interactions the time-integrated Berry curvature ${\cal W}_{k}^{-}$ shows a discontinuity at the light lines at $k_{d}a=\pm\pi/2$ (see Fig. 7d). This again constraints the width of the Gaussian wave packet in momentum space, which must satisfy $w_{k}\ll\pi/a$ in order to achieve dispersionless pumping. Here, the role of the flatness of ${\cal W}_{k}^{-}$ is made particularly clear: since the light lines are closer to $k=0$ than in the case of low-lying two-level systems, the curvature of ${\cal W}_{k}^{-}$ within the region of the Gaussian wave packet is larger, which in turn reduces the fidelity of the pumping process. Finally, we note that in the waveguide case the pumping occurs in the opposite direction compared with the two previous cases studied. This is due to the sign of ${\cal W}_{k}^{-}$ which is negative around $k=0$, while the Chern number of the band, i.e., the integral of ${\cal W}_{k}^{-}$ over the full FBZ, is the same as in all systems, namely $C_{-}=+1$. Note, that in principle the direction of pumping is arbitrary and depends on the orientation of the path around the singularity, i.e. the excitation moves in opposite direction by reversing the path shown in Fig. 5-7c.
@@ 5 Robustness against dissipation and disorder
Up until now, we have shown that the topological pumping of a photon in a system with long-range interactions is possible. However, these quantum optical systems also suffer from dissipation, e.g. radiative decay of the photon into the environment, as introduced in equation (7). Moreover, experimental systems have a degree of disorder due to, for example, imperfect trapping of the atoms, which leads to disorder in the hopping rates. In this section, we will analyze the effect of these two potentially detrimental sources on the topological photon pumping protocols.
Figure 8: Collective dissipation. Survival probability for the photon as a function of time during the pumping cycle for a chain with $a/\lambda_{0}=0.4$ (red solid line) and $a/\lambda_{0}=0.7$ (blue dash-dotted line). For $a/\lambda_{0}=0.4$, $|f(k)|^{2}$ is nonzero only in a region of the FBZ where $\Gamma_{k}=0$, leading to a survival probability $P_{\mathrm{sur}}(t)=1$ for the full cycle. The same wave packet for $a/\lambda_{0}=0.7$ has a large decay rate, apparent from the quick decay of the survival probability. Note, that in both cases we choose an angle $\theta_{d}$ such that, as discussed previously, $J_{2}=0$ and the sublattice symmetry is approximately conserved.
@@ 5.1 Dissipation
Dissipation has a different character for the three platforms we have studied here. In all three, however, one can encode it in its most general form as the action of the dissipator
@@
i.e., as in the case of the dipole-dipole interactions (23), it depends strongly on the ratio between the interatomic distance and the transition wavelength, the precise external geometry of the system and the orientation of the transition dipole moments. The (collective) jump operators and rates can be found by diagonalizing this matrix, such that $\Gamma_{k}=\sum_{ij}M_{ki}\Gamma_{ij}M^{*}_{jk}$ and $\hat{J}_{k}=\sum_{j}M_{kj}\hat{\sigma}_{j}$. Here one can see that the jump operators are then in general superpositions of all single atom spin operators $\hat{\sigma}_{j}=\ket{g}_{j}\!\!\bra{e}$ |, with rates that can be larger (superradiant) or smaller (subradiant) than the single atom decay rate. In the regime we consider here, a small number of superradiant states with $N\gamma>\Gamma_{k}$ are present. The region of the FBZ where these superradiant states appear, however, depends strongly on the specific value of the ratio between nearest neighbor distance and $\lambda_{0}$. In particular, due to the large sublattice energy shifts $\Delta$ that are present for most of the pumping cycle, the two sublattices are effectively decoupled dynamically, which means that the reduced distance that determines the dissipation properties is $a/\lambda_{0}$. As can be observed in Fig. 8, the superradiant rates can be found at the center and the border of the FBZ for $a/\lambda_{0}<1/2$ and $a/\lambda_{0}>1/2$, respectively. This in turn means that, as shown in Fig. 8, choosing a value $a/\lambda_{0}<1/2$, one can create a wave packet centered at $k=\pm\pi/a$ with support exclusively in the subradiant region, which in turn leads to a survival probability of the photon $P_{\mathrm{sur}}$ equal to one for all experimentally relevant times. For comparison, the same wave packet created on a chain where $a/\lambda_{0}>1/2$ will decay fast (in the example we show in Fig. 8 with $a/\lambda_{0}=0.7$, $\Gamma_{\mathrm{eff}}=1.6183\gamma$).
@@
Figure 9: Robustness against disorder. Fidelity after one pumping cycle $\mathcal{F}_{\sigma}(T)$ between the disordered and non-disordered states as a function of the time-integrated disorder in the extended hopping rates $\sigma(\delta\mathcal{J})/\delta{\mathcal{J}}$. The insets show the disordered paths in parameter space for the three systems with a time-integrated disorder of 10%. The unperturbed path in parameter space is shown by the red line, while the blue shaded area shows the propagated disorder given a Gaussian disorder in each emitter position in the chain.
@@ 5.2 Disorder
as a common measure of disorder for the three systems. Hence, in order to get a fair comparison, we choose the same value of the time-integrated disorder $\sigma(\delta\mathcal{J})/\delta\mathcal{J}$ for all systems. Note, however, that as a result of propagation of disorder the Rydberg system with short range interactions requires a larger disorder in the position compared to the two other systems to reach a equivalent disorder in the path.
To measure the effect of disorder, we calculate the fidelity $\mathcal{F}_{\sigma}(T)=|\langle\Psi_{T}^{\sigma}|\Psi_{T}^{0}\rangle|^{2}$, where $|\Psi_{T}^{\sigma}\rangle$ is the disordered state averaged over a number of realizations of random Gaussian disorder in the emitter positions and $|\Psi_{T}^{0}\rangle$ is the non-disordered state after one cycle. Figure 9 shows this fidelity for the three different systems as a function of the time-integrated disorder $\sigma(\delta\mathcal{J})/\delta\mathcal{J}$ up to 10%. We can observe that the pumping is indeed very robust in all three cases, with the atoms in low-lying states being the most affected, likely due to the slight breaking of the sublattice symmetry present in this case.
@@ 6 Conclusions and outlook
This study establishes for the first time a framework for topological photon pumping in quantum optical systems with long-range interactions. This framework allows for the controlled and robust transport of photonic excitations over a one-dimensional emitter chain. Notably, the atoms are only displaced over short distances within a single unit cell, while the excitation is transported with high fidelity over many lattice sites. We discuss its realization in different experimental platforms, such as Rydberg atoms, in a dense atomic lattice gas, and atoms coupled to waveguides. For each system we establish a counter-intuitive protocol that displaces only half of the atoms, but allows to transport the photonic excitation with a fidelity reaching 99.9%. Our results underscore that this scheme not only shifts the center-of-mass of the excitation, but also the maintains the shape of the wave packet (dispersionless) and we show that it is robust against both local disorder and dissipation.
@@ Appendix A Generalized expression for $\Delta x$
We consider the initial state of the dynamics to be a superposition of Bloch waves, i.e. |$\ket{\Psi_{0}(x)}=\sum_{k}f(k)\ket{\psi_{nk}(x)}$, where $f(k)$ is the quasi-momentum distribution normalized such that
$\int_{\mathrm{FBZ}}|f(k)|^{2}\mathrm{d}k=\frac{2\pi}{a}.$
Initially, expectation value of the position is given by $\langle\hat{x}\rangle_{t}\equiv\sum_{x}\bra{\Psi_{t}(x)}\hat{x}\ket{\Psi_{t}(x)}$. Expressing this in terms of the superposition of Bloch waves
$\langle\hat{x}\rangle_{t}=\sum_{k,k^{\prime},x}f(k)f^{*}(k^{\prime})\bra{u_{nk^{\prime}}(t)}xe^{ix(k-k^{\prime})}\ket{u_{nk}(t)}\!.$
Now, transforming the sum over $k$ to an integral
$\begin{split}\langle\hat{x}\rangle_{t}=\frac{1}{\sqrt{2\pi}}&\sum_{k^{\prime},x}f^{*}(k^{\prime})\bra{u_{nk^{\prime}}(t)}e^{-ik^{\prime}x}\\ &\times\int_{\text{FBZ}}\mathrm{d}kf(k)e^{ikx}x\ket{u_{nk}(t)}.\end{split}$
@@
$\begin{split}&\langle\hat{x}\rangle_{t}=\frac{i}{\sqrt{2\pi}}\sum_{k^{\prime},x^{\prime}}f^{*}(k^{\prime})\bra{u_{nk^{\prime}}(t)}e^{-ik^{\prime}x}\\ &\times\int_{\text{FBZ}}\!\!\!\mathrm{d}ke^{ikx}[\partial_{k}f(k)\ket{u_{nk}(t)}+f(k)\partial_{k}\ket{u_{nk}(t)}].\end{split}$
Performing the sum over real-space, $\sum_{x}e^{i(k-k^{\prime})x}=a\delta_{kk^{\prime}}$, and transforming the sum over $k^{\prime}$ into an integral we obtain
$\begin{split}\langle\hat{x}\rangle_{t}&=\frac{ia}{2\pi}\!\int_{\text{FBZ}}\mathrm{d}k\bigg{[}f(k)\partial_{k}f^{*}(k)\\ &+|f(k)|^{2}\bra{u_{nk}(t)}\partial_{k}u_{nk}(t)\rangle\!\bigg{]}\!.\end{split}$
@@
$\begin{split}\Delta x=&\langle\hat{x}\rangle_{T}-\langle\hat{x}\rangle_{0}=\int_{0}^{T}\mathrm{d}t\,\partial_{t}\langle\hat{x}\rangle_{t}\\ =&\frac{a}{2\pi}\int_{0}^{T}\mathrm{d}t\int_{\text{FBZ}}\mathrm{d}k|f(k)|^{2}\partial_{t}A_{nk}(t),\end{split}$
where we have introduced the Berry connection $A_{nk}(t)=i\bra{u_{nk}(t)}\partial_{k}u_{nk}(t)\rangle$. The integral may now be expressed in term of the Berry curvature of the band $\Omega_{tk}=\partial_{t}A_{nk}-\partial_{k}A_{nt}=-2\text{Im}\{\langle\partial_{t}u_{nk}(t)|\partial_{k}u_{nk}(t)\rangle\}$ as
$\begin{split}&\Delta x=\frac{a}{2\pi}\int_{0}^{T}\mathrm{d}t\int_{\mathrm{FBZ}}|f(k)|^{2}\Omega_{tk}\mathrm{d}k\\ &+\frac{ia}{2\pi}\int_{0}^{T}\mathrm{d}t\int_{\mathrm{FBZ}}\mathrm{d}k|f(k)|^{2}\partial_{k}\bra{u_{nk}(t)}\partial_{t}u_{nk}(t)\rangle.\end{split}$
Using partial integration on the last integral over $k$ and using the periodicity of the function $\bra{u_{nk}(t)}\partial_{t}u_{nk}(t)\rangle$ in the FBZ, this term vanishes and we end up with the final expression for the displacement of the center of mass
$\Delta x=\frac{a}{2\pi}\int_{0}^{T}\mathrm{d}t\int_{\mathrm{FBZ}}|f(k)|^{2}\Omega_{tk}\mathrm{d}k,$

### v3

@@ 4.2.1 Photon pumping on a two-level atom chain
Figure 7: Photon pumping on a waveguide system. a: Dependence of the extended hopping parameters $\bar{J}^{\prime}$ and $\bar{J}$ with the angle $\phi^{\prime}_{1}$ (all parameters are normalized to the maximal value $\bar{J}_{\text{max}}\equiv|\text{max}\{\bar{J}(t)\}|$ during the cycle). The gray shaded area represents $\delta\bar{J}<0$, where the winding number of the extended SSH model is $\nu=1$ (TP: topological phase). b: Variation of the angle $\phi^{\prime}_{1}(t)$ during the pumping cycle. Importantly, twice during the cycle the extended hopping parameters coincide, such that $\delta\bar{J}=0$. c: We vary the energy offset as $\Delta(t)=\Delta_{\mathrm{max}}\sin(2\pi t/T+\pi/2)$ such that the path in parameter space encircles the topological degeneracy point ensuring the displacement of the center of mass after a completed cycle. d: The left column shows the time-integrated Berry curvature of the lower band ${\cal W}_{k}^{-}$ and quasi-momentum distribution density $|f(k)|^{2}$ (in arbitrary units). The right column shows the spatial probability density of the wave packet in real space (on the $B$ sublattice), initially (dashed line) and after its numerical evolution for ten pumping cycles (solid line). The width of the Gaussian wave packet in momentum space $w_{k}$ (the same as in Figs. 5 and 6) is doubled from the upper to the lower row, leaving the displacement of the photon’s center of mass $\Delta x=10\cdot 0.919a$ unchanged, but leading to an increase of the dispersion, such that the fidelity per cycle ${\cal F}(T)=0.998$ and after ten cycles ${\cal F}(10T)=0.962$ decreases to 0.987 and 0.540, respectively. Here, $T=150/\bar{J}_{\text{max}}$.
@@ 4.3 All-to-all hopping: Waveguide system
Finally, we discuss a chain of emitters coupled to a waveguide. This setup can be realized, for instance, with cold atoms trapped in the vicinity of a tapered optical fiber [57, 75, 76]. In this system, the coherent interactions between atoms are mediated by the fundamental guided modes of the waveguide. Note, that we limit the discussion in this paper to cases where the coupling efficiency of the emitters to the guided modes is almost $100\%$, as it can be achieved, for example, with artificial atoms in semiconductors (quantum dots) [77, 78]. In general, however, one needs to consider the coupling through the radiation or unguided modes. In the case of atoms placed around a cylindrical nanofiber (see Fig. 4d), the interactions between atoms separated by an angle $\phi_{ij}=\phi_{j}-\phi_{i}$ and a distance $z_{ij}=z_{j}-z_{i}$ are given by
@@
where $\gamma=\frac{2|d|^{2}\omega_{0}\beta^{\prime}}{\hbar\epsilon_{0}}|e_{r}|^{2}$ is the rate of a photon to be emitted into the waveguide for a single atom. Moreover, we consider in particular the situation where the atomic cells are positioned as a helix around the waveguide as shown in Fig. 4d  [79]. The hopping parameters in this case are given by
@@
where $\phi_{1}$ and $\phi_{1}^{\prime}$ are the angle between atoms in neighboring cells of the same sublattice, and between atoms of different sublattices within the same cell, respectively. In order to ensure that the chiral symmetry is not broken, we fix the value of $a$ to $a=\pi/\beta$. We also fix the values of $b=a/2=\pi/(2\beta)$ and $\phi_{1}=\pi/2$, such that $J_{2p}=0$ and the remaining hopping parameters read
@@ 4.3.1 Photon pumping on a waveguide system
As shown in Fig. 7a, the transition from a topological to a non-topological phase occurs in this system when the angle $\phi_{1}^{\prime}=\pi/4$, since at this point $J_{2p-1}=J^{\prime}_{2p-1}$ for all values of $p$. A pumping scheme can then be established by adiabatically rotating sublattice $B$ by an angle $\phi_{1}^{\prime}(t)$ which after a time $T$ returns to the initial configuration (Figs. 7b and c). As it was discussed for atoms in free space (sec Sec. 4.2), due to the long-ranged character of the couplings the time-integrated Berry curvature ${\cal W}_{k}^{-}$ shows a discontinuity at the light lines at $k_{d}a=\pm\pi/2$ (see Fig. 7d). This again constrains the width of the Gaussian wave packet in momentum space, which must satisfy $w_{k}\ll\pi/a$ in order to achieve dispersionless pumping. Here, the role of the flatness of ${\cal W}_{k}^{-}$ is made particularly clear: since the light lines are closer to $k=0$ than in the case of low-lying two-level systems, the curvature of ${\cal W}_{k}^{-}$ within the region of the Gaussian wave packet is larger, which in turn reduces the fidelity of the pumping process. Finally, we note that in the waveguide case the pumping occurs in the opposite direction compared with the two previous cases studied. This is due to the sign of ${\cal W}_{k}^{-}$ which is negative around $k=0$, while the Chern number of the band, i.e., the integral of ${\cal W}_{k}^{-}$ over the full FBZ, is the same as in all systems, namely $C_{-}=+1$. Note, that in principle the direction of pumping is arbitrary and depends on the orientation of the path around the singularity, i.e. the excitation moves in opposite direction by reversing the path shown in Fig. 5-7c.
Figure 8: Collective dissipation. a: Survival probability for the photon as a function of time during the pumping cycle for a chain with $a/\lambda_{0}=0.4$ (red solid line) and $a/\lambda_{0}=0.7$ (blue dash-dotted line). b: For $a/\lambda_{0}=0.4$, $|f(k)|^{2}$ is nonzero only in a region of the FBZ where $\Gamma_{k}=0$, leading to a survival probability $P_{\mathrm{sur}}(t)=1$ for the full cycle. c: The same wave packet for $a/\lambda_{0}=0.7$ has a large decay rate, apparent from the quick decay of the survival probability. Note, that in both cases we choose an angle $\theta_{d}$ such that, as discussed previously, $J_{2}=0$ and the sublattice symmetry is approximately conserved.
@@ 5 Robustness against dissipation and disorder
Up until now, we have shown that the topological pumping of a photon in a system with long-range couplings is possible. However, these quantum optical systems also suffer from dissipation, e.g. radiative decay of the photon into the environment, as introduced in equation (7). Moreover, experimental systems have a degree of disorder due to, for example, imperfect trapping of the atoms, which leads to disorder in the hopping parameters. In this section, we will analyze the effect of these two potentially detrimental sources on the topological photon pumping protocols.
@@ 5.1 Dissipation
Dissipation has a different character for the three platforms we have studied here. In all three, however, one can encode it in a general form as the action of the dissipator
@@
i.e., as in the case of the dipole-dipole interactions (23), it depends strongly on the ratio between the interatomic distance and the transition wavelength, the precise external geometry of the system and the orientation of the transition dipole moments [73]. The (collective) jump operators and rates can be found by diagonalizing this matrix, such that $\Gamma_{k}=\sum_{ij}M_{ki}\Gamma_{ij}M^{*}_{jk}$ and $\hat{J}_{k}=\sum_{j}M_{kj}\hat{\sigma}_{j}$. Here one can see that the jump operators are then in general superpositions of all single atom spin operators $\hat{\sigma}_{j}=\ket{g}_{j}\!\!\bra{e}$ |, with rates that can be larger (superradiant) or smaller (subradiant) than the single atom decay rate. In the regime we consider here, a small number of superradiant states with $N\gamma>\Gamma_{k}$ are present. The region of the FBZ where these superradiant states appear, however, depends strongly on the specific value of the ratio between nearest neighbor distance and $\lambda_{0}$. In particular, due to the large sublattice energy shifts $\Delta$ that are present for most of the pumping cycle, the two sublattices are effectively decoupled dynamically, which means that the reduced distance that determines the dissipation properties is $a/\lambda_{0}$. As can be observed in Figs. 8b and c, the superradiant rates can be found at the center and the border of the FBZ for $a/\lambda_{0}<1/2$ and $a/\lambda_{0}>1/2$, respectively. This in turn means that, as shown in Fig. 8b, choosing a value $a/\lambda_{0}<1/2$, one can create a wave packet centered at $k=\pm\pi/a$ with support exclusively in the subradiant region, which in turn leads to a survival probability of the photon $P_{\mathrm{sur}}$ equal to one for all experimentally relevant times. For comparison, the same wave packet created on a chain where $a/\lambda_{0}>1/2$ will decay fast (in the example we show in Fig. 8a with $a/\lambda_{0}=0.7$, $\Gamma_{\mathrm{eff}}=1.6183\gamma$).
@@
Figure 9: Robustness against disorder. Fidelity after one pumping cycle $\mathcal{F}_{\sigma}(T)$ between the disordered and non-disordered states as a function of the time-integrated disorder in the extended hopping parameters $\sigma(\delta\mathcal{J})/\delta{\mathcal{J}}$. The insets show the disordered paths in parameter space for the three systems with a time-integrated disorder of 10%. The unperturbed path in parameter space is shown by the red line, while the blue shaded area shows the propagated disorder given a Gaussian disorder in each emitter position in the chain.
@@ 5.2 Disorder
as a common measure of disorder for the three systems. Hence, in order to get a fair comparison, we choose the same value of the time-integrated disorder $\sigma(\delta\mathcal{J})/\delta\mathcal{J}$ for all systems. Note, however, that as a result of propagation of disorder the Rydberg system with short range couplings requires a larger disorder in the position compared to the two other systems to reach a equivalent disorder in the path.
To measure the effect of disorder, we calculate the fidelity
$\mathcal{F}_{\sigma}(T)=\left(\mathrm{Tr}\sqrt{\sqrt{\rho_{T}^{0}}\rho_{T}^{\sigma}\sqrt{\rho_{T}^{0}}}\right)^{2},$
where $\rho_{T}^{\sigma}$ is the disordered state averaged over a number of realizations of random Gaussian disorder in the emitter positions and $\rho_{T}^{0}=|\Psi_{T}^{0}\rangle\langle\Psi_{T}^{0}|$ is the non-disordered state after one cycle. Figure 9 shows this fidelity for the three different systems as a function of the time-integrated disorder $\sigma(\delta\mathcal{J})/\delta\mathcal{J}$ up to 10%. We can observe that the pumping is indeed very robust in all three cases, with the atoms in low-lying states being the most affected, likely due to the slight breaking of the sublattice symmetry present in this case, i.e. the presence of $J_{2p}\neq 0$ terms, which effectively double the amount of couplings affected by the disorder.
@@ 6 Conclusions and outlook
This study establishes for the first time a framework for topological photon pumping in quantum optical systems with long-range couplings. This framework allows for the controlled and robust transport of photonic excitations over a one-dimensional emitter chain. Notably, the atoms are only displaced over short distances within a single unit cell, while the excitation is transported with high fidelity over many lattice sites. We discuss its realization in different experimental platforms, such as Rydberg atoms, in a dense atomic lattice gas, and atoms coupled to waveguides. For each system we establish a counter-intuitive protocol that displaces only half of the atoms, but allows to transport the photonic excitation with a fidelity reaching 99.9%. Our results underscore that this scheme not only shifts the center-of-mass of the excitation, but also maintains the shape of the wave packet (dispersionless) and we show that it is robust against both local disorder and dissipation.
@@
While this manuscript was under review, we became aware of an experimental realization of a charge pump with Rydberg excitations [84].
@@ Appendix A Generalized expression for $\Delta x$
We consider the state of the system throughout the dynamics to be a superposition of Bloch waves, i.e.
|$\ket{\Psi_{t}(x)}=\frac{1}{\sqrt{N}}\sum_{k=-\pi/a}^{\pi/a}f(k)\ket{\psi_{nk}(x)},$
where $N$ is the number of unit cells, $f(k)$ is the quasi-momentum distribution normalized such that
$\int_{\mathrm{FBZ}}|f(k)|^{2}\mathrm{d}k=\frac{2\pi}{a},$
and
|$\ket{\psi_{nk}(x)}=\frac{1}{\sqrt{N}}e^{ikx}\ket{x}\otimes\ket{u_{nk}(t)}.$
In the limit of an infinite system, i.e., $N\to\infty$, $k$ becomes a continuous variable. Thus, we can substitute the sum over $k$, $\sum_{k=-\pi/a}^{\pi/a}$, by the corresponding integral, $\frac{Na}{2\pi}\int_{\mathrm{FBZ}}\mathrm{d}k$, such that the state reads
|$\ket{\Psi_{t}(x)}=\frac{a}{2\pi}\int_{\text{FBZ}}\mathrm{d}kf(k)e^{ikx}\ket{x}\otimes\ket{u_{nk}(t)}.$
The expectation value of the position as a function of time is given by
$\begin{split}\langle\hat{x}\rangle_{t}&=\sum_{x=-\infty}^{\infty}\bra{\Psi_{t}(x)}\hat{x}\ket{\Psi_{t}(x)}\\ &=\left(\frac{a}{2\pi}\right)^{2}\sum_{x}\int_{\text{FBZ}}\int_{\text{FBZ}}\mathrm{d}k\mathrm{d}k^{\prime}f(k)f^{*}(k^{\prime})\\ &\times\bra{u_{nk^{\prime}}(t)}xe^{ix(k-k^{\prime})}\ket{u_{nk}(t)}.\end{split}$
@@
$\begin{split}&\langle\hat{x}\rangle_{t}=i\left(\frac{a}{2\pi}\right)^{2}\sum_{x}\int_{\text{FBZ}}\mathrm{d}k^{\prime}f^{*}(k^{\prime})\bra{u_{nk^{\prime}}(t)}e^{-ik^{\prime}x}\\ &\times\!\int_{\text{FBZ}}\!\!\!\mathrm{d}ke^{ikx}\{\left[\partial_{k}f(k)\right]\ket{u_{nk}(t)}+f(k)\partial_{k}\ket{u_{nk}(t)}\}.\end{split}$
Using the Fourier decomposition of the Dirac delta function $\sum_{x=-\infty}^{\infty}e^{i(k-k^{\prime})x}=\frac{2\pi}{a}\delta(k-k^{\prime})$, we obtain
$\begin{split}\langle\hat{x}\rangle_{t}&=\frac{ia}{2\pi}\!\int_{\text{FBZ}}\mathrm{d}k\bigg{[}f^{*}(k)\partial_{k}f(k)\\ &+|f(k)|^{2}\bra{u_{nk}(t)}\partial_{k}u_{nk}(t)\rangle\!\bigg{]}\!.\end{split}$
@@
$\begin{split}\Delta x=&\langle\hat{x}\rangle_{T}-\langle\hat{x}\rangle_{0}=\\ =&\frac{a}{2\pi}\int_{\text{FBZ}}\mathrm{d}k|f(k)|^{2}\left[A_{nk}(T)-A_{nk}(0)\right],\end{split}$
where we have introduced the Berry connection $A_{nk}(t)=i\bra{u_{nk}(t)}\partial_{k}u_{nk}(t)\rangle$. Considering that, after a full adiabatic cycle, the time-evolved eigenstates only pick up the so-called geometric or Berry phase, i.e. |$\ket{u_{nk}(T)}=e^{i\gamma(k)}\ket{u_{nk}(0)}$, one can easily find that
$\begin{split}\Delta x&=-\frac{a}{2\pi}\int_{\mathrm{FBZ}}|f(k)|^{2}\partial_{k}\gamma(k)\mathrm{d}k\\ &=\frac{a}{2\pi}\int_{0}^{T}\mathrm{d}t\int_{\mathrm{FBZ}}|f(k)|^{2}\Omega_{tk}\mathrm{d}k,\end{split}$